# Molecular Generative Adversarial Network with Multi-Property Optimization

### v1

@@ Abstract
Deep generative models, such as generative adversarial networks (GANs), have been employed for $de~{}novo$ molecular generation in drug discovery. Most prior studies have utilized reinforcement learning (RL) algorithms, particularly Monte Carlo tree search (MCTS), to handle the discrete nature of molecular representations in GANs. However, due to the inherent instability in training GANs and RL models, along with the high computational cost associated with MCTS sampling, MCTS RL-based GANs struggle to scale to large chemical databases. To tackle these challenges, this study introduces a novel GAN based on actor-critic RL with instant and global rewards, called InstGAN, to generate molecules at the token-level with multi-property optimization. Furthermore, maximized information entropy is leveraged to alleviate the mode collapse. The experimental results demonstrate that InstGAN outperforms other baselines, achieves comparable performance to state-of-the-art models, and efficiently generates molecules with multi-property optimization. The source code will be released upon acceptance of the paper.
Machine Learning, ICML
@@ 1 Introduction
Molecular graphs (Shi et al., 2020) and simplified molecular input line entry systems (SMILES) strings (Weininger, 1988) constitute the two primary representations of molecules in deep generative models. However, generating molecules with desired chemical properties using such discrete representations is a non-trivial task. Most prior studies related to generative adversarial networks (GANs) (Yu et al., 2017; Guimaraes et al., 2017; De Cao & Kipf, 2018) typically update the generator by integrating the output probability of the discriminator with the chemical properties of generated molecules as a reward for reinforcement learning (RL), following the REINFORCE algorithm (Williams, 1992). Due to the inability of GANs to calculate rewards for partially generated molecules, Monte Carlo tree search (MCTS) is frequently utilized for sampling and completing molecules (Li et al., 2022). Unfortunately, the integration of RL algorithms with GANs further exacerbates the instability of the training process. Stabilizing MCTS demands a substantial number of samples, significantly lengthening the process (Li & Yamanishi, 2023).
@@
Furthermore, most aforementioned studies on $de~{}novo$ molecular generation have focused on optimizing a single chemical property. However, in practical applications, it is often desirable to generate molecules that satisfy multiple chemical property constraints. In contrast to the former, multi-property optimization is highly complicated and challenging to achieve in nature. This is because the multi-property optimization task entails not only learning the semantic and syntactic rules of molecules to generate valid molecules from scratch but also finding pathways to optimize the distribution of chemical properties during the process (Barshatski et al., 2021). For example, molecules exhibiting both drug-likeness and dopamine receptor (DRD2) activity represent only $1.6\%$ of the generated molecules (Jin et al., 2019). Therefore, employing deep generative models for $de~{}novo$ molecular generation with multi-property optimization holds significance to the drug discovery industry (Barshatski & Radinsky, 2021).
@@
Inspired by the previous studies in (De Cao & Kipf, 2018; Tang et al., 2023), this study introduces a novel GAN based on actor-critic RL with instant rewards (IR) and global rewards (GR), called InstGAN, to generate molecules at the token-level with multi-property optimization. Specifically, the generator is constructed using a long short-term memory (LSTM) that generates SMILES strings in an autoregressive manner. The discriminator quantifies each token based on SMILES substrings produced by the generator. To enhance the ability of the discriminator to quantize tokens, a bidirectional LSTM (Bi-LSTM) is chosen as the discriminator. Additionally, multiple property prediction networks with the same structure as the discriminator predict the corresponding property scores for each token as well. Subsequently, the scores of discriminator and property prediction networks, along with their global-level scores, serve as rewards for RL. To expedite the training process and facilitate its application to extensive chemical databases, an IR calculation based on actor-critic RL is proposed to update the generator. Furthermore, the maximized information entropy (MIE) is included in the loss function for generator updates to mitigate mode collapse and enhance molecular diversity. The main contributions are as follows:
@@ 2.1 Variational autoencoder (VAE)-based Models
Two VAE variants, Character-VAE and Grammar-VAE (Kusner et al., 2017), integrate parse trees with VAEs to facilitate the generation of syntactically valid molecules. However, due to ignoring the semantic information of the molecular representations, the correlation between the training set and the generated molecules cannot be guaranteed. In contrast, Syntax-VAE (Dai et al., 2018) ensures that the generated molecules are both syntactically valid and semantically meaningful. JT-VAE (Jin et al., 2018) adopts a two-step VAE approach, first generating tree-structured scaffolds based on chemical substructures and then combining these outputs into complete molecules using a graph message-passing network (Dai et al., 2016; Gilmer et al., 2017). Nonetheless, a major limitation of VAEs is the typically limited size of the latent space, which may restrict the capacity to produce highly realistic molecules.
@@ 2.2 Flow-based Models
GraphAF (Shi et al., 2020) utilizes a flow-based autoregressive model to generate molecular graphs. GraphDF (Luo et al., 2021) incorporates invertible modulo shift transforms to connect discrete latent variables with graph nodes and edges, resulting in the generation of molecular graphs. MoFlow (Zang & Wang, 2020) adopts a Glow-based model (Kingma & Dhariwal, 2018) to generate chemical bonds as graph edges and employs a graph conditional flow to subsequently generate atoms as graph nodes, followed by posthoc validity correction. GraphCNF (Lippe & Gavves, 2021) falls under the category of flow-based models, commonly applied in various data domains. In molecular graph generation, GraphCNF leverages flow-based techniques to address the unique challenges associated with generating molecular graphs. However, flow-based models generally exhibit two notable limitations: the computation of the Jacobian matrix is time-consuming, often requiring approximations, and the necessity for network invertibility imposes constraints on their representational capacity (Zhang et al., 2021).
@@ 2.3 Diffusion-based Models
Recently, diffusion models have found application in the domain of molecular generation, where they are being utilized to tackle the intricate challenges associated with generating molecular structures that adhere to specific chemical and property constraints. DiGress (Vignac et al., 2023) iteratively refines noisy graphs by adding or removing edges and adjusting categories, which results in the generation of molecular graphs with node and edge attributes suitable for classification. GDSS (Jo et al., 2022) and D2L-OMP (Guo et al., 2023) are the two SOTA models in the realm of molecular generation. GDSS skillfully captures the joint distribution of molecular nodes and edges, generating molecular replicas closely aligned with the training distribution. D2L-OMP generates a hybrid Gaussian distribution by employing diffusion on two structural levels: molecules and molecular fragments. This hybrid Gaussian distribution is then utilized in the reverse denoising process.
Figure 1: The architectural overview of InstGAN comprises three main substructures. (a) The generator, featuring an LSTM, produces tokens in an autoregressive manner at each time step. MIE is employed to enhance the likelihood of sampling different tokens, smoothing the output probability distribution for the generator. (b) The discriminator, a Bi-LSTM, scores generated tokens based on both forward and backward at each time step, enabling semantic and syntactic discrimination at the token-level. Higher probabilities are assigned to likely generated tokens, while those with errors receive lower probabilities. The average of all-token probabilities assesses overall generation quality from a global perspective. (c) Multiple pretrained chemical property prediction networks, labeled Critic $1$ to Critic $N$, have the same structure as the discriminator. They predict various chemical properties for each token of a generated SMILES string. Similarly, the sum of stepwise scores is averaged to provide a global property score for the entire SMILES string. Finally, scores from the discriminator and critics serve as rewards in the actor-critic RL algorithm, co-updating the generator via the policy gradients.
@@ 2.4 GAN-based Models
SeqGAN (Yu et al., 2017) pioneered the incorporation of MCTS-based RL into GAN architecture, specifically designed to handle discrete text. This innovation has inspired various studies on molecular generation using GANs. MolGAN (De Cao & Kipf, 2018) introduces an implicit, likelihood-free discrete GAN for generating small molecular graphs. However, MolGAN faces an overfitting problem, leading to less than 5% uniqueness in the generated molecules. ORGAN (Guimaraes et al., 2017) integrates domain-specific knowledge as rewards for generating SMILES strings through the MCTS-based RL algorithm. TransORGAN (Li et al., 2022) leverages a transformer architecture to capture semantic information and employs variant SMILES technique for syntax rule learning. SpotGAN (Li & Yamanishi, 2023) adopts a first-decoder-then-encoder transformer model for generating SMILES strings from a given scaffold. However, the use of MCTS-based RL in GANs often demands a substantial number of samples for training stability, making it impractical for extensive chemical databases. EarlGAN (Tang et al., 2023), while capable of generating valid molecules on large chemical databases, lacks the ability to optimize chemical properties, particularly multiple properties simultaneously. This study introduces an actor-critic RL-driven GAN that employs IR and GR to enhance the efficiency of learning semantic and syntactic rules in SMILES strings. Furthermore, our other goal is to identify pathways to optimize the generation of molecules across a broad spectrum of chemical properties.
@@ 3 Model
Figure 1 illustrates the overall architecture of InstGAN and highlights the three key substructures (i.e., the generator, discriminator, and multiple chemical property prediction networks) that are crucial for generating molecules with multi-property optimization from scratch.
Formally, let $G_{\bm{\theta}}$ and $D_{\bm{\phi}}$ represent the generator and discriminator of InstGAN with parameters $\bm{\theta}$ and $\bm{\phi}$, respectively. ${\bm{S}}_{1:T}=[s_{1},\cdots,s_{T}]$ denotes a SMILES string with a sequence length of $T$. Then, the min-max optimization procedure (Goodfellow et al., 2014) is as follows:
$\min_{\bm{\theta}}\max_{\bm{\phi}}V(G_{\bm{\theta}},D_{\bm{\phi}})=$ $\mathbb{E}_{{\bm{S}}\sim p_{r}({\bm{S}})}[\log D_{\bm{\phi}}({\bm{S}})]+$
$\mathbb{E}_{{\bm{S}}\sim p_{\bm{\theta}}({\bm{S}})}[\log(1-D_{\bm{\phi}}({\bm{S}}))],$ (1)
where $p_{r}(\cdot)$ and $p_{\bm{\theta}}(\cdot)$ represent the distributions of training SMILES strings and generated sequences, respectively. Following a similar approach to NLP methods (de Masson d’Autume et al., 2019; Fedus et al., 2018), InstGAN utilizes an autoregressive generator and discriminator. Notably, this design allows for the allocation of dense rewards at the token-level. See Appendix A for more details on InstGAN’s generator and discriminator.
@@ 3.1 Token-Level Critics
Unlike graph-based approaches, molecular generative models relying on SMILES strings often face challenges in ensuring high validity due to the intricacies of checking valence during the autoregressive generation process. Typically, invalid SMILES strings may arise from mismatched tokens, requiring the removal or replacement of other suitable tokens to restore validity. Therefore, a crucial aspect of addressing this issue lies in the meticulous assessment of SMILES strings at the token-level.
@@
In InstGAN, the generator employs an autoregressive process to produce SMILES strings, serving as inputs for the discriminator. Diverging from traditional discriminators, InstGAN performs token-level discrimination for each generated token. Specifically, with a Bi-LSTM, let $\overrightarrow{\bm{S}}_{1:t}$ and $\overleftarrow{\bm{S}}_{t:T}$ represent the forward and backward SMILES substrings, respectively. The discriminator calculates the probability $r^{D}_{t}$ that it deems the $t$-th token $\widetilde{s}_{t}$ of the SMILES string as true. The calculation is formulated as follows:
@@
To assess the overall validity of the entire SMILES string, we calculate the global discriminator score as follows:
@@
The chemical property prediction networks are replicated from the discriminator and share the same structure. In the pretraining phase, the entire SMILES string serves as input, and the networks utilize the corresponding property values of the complete SMILES string as labels for the chemical properties of its tokens at each step. In the training phase of InstGAN, these networks calculate both instant and global chemical properties of the generated SMILES string. Collaborating with the discriminator, they contribute to updating the generator using actor-critic RL-based rewards.
@@ 3.2 Instant and Global Reward Calculation
Following the actor-critic RL algorithm, the generator functions as the actor network responsible for action selection, while the discriminator, along with the property prediction networks, serves as the critic for reward calculation. However, in contrast to the traditional actor-critic RL algorithm (Konda & Tsitsiklis, 1999), which calculates the reward at the last time step $T$, we compute the reward for each token of the discriminator as follows:
@@
where $W^{D}$ represents the weight assigned to the discriminator’s GR. Similarly, the reward of the property prediction networks can be calculated and denoted as $R^{C_{i}}_{t}$, $i\in[1,2,\cdots,N]$, and $N$ indicates the number of chemical properties to be optimized. Subsequently, the overall reward for chemical properties is represented as follows:
$R^{C}_{t}=\sum_{i=1}^{N}W_{C_{i}}R^{C_{i}}_{t}.$ (5)
Here, $W_{C_{i}}$ denotes the weight assigned to the $i$-th chemical property prediction network. Finally, the total reward $R_{t}$ at time step $t$ can be calculated as
@@
where $\lambda$ represents a hyperparameter that balances the trade-off between the GAN and RL components.
@@
2:  Initialization: generator: $G_{\bm{\theta}}$, discriminator: $D_{\bm{\phi}}$, property prediction networks: $C^{i}_{{\bm{\varphi}}_{i}},i\in[1,\cdots,N]$
@@
4:  // Pretrain the discriminator
[…]
@@
8:  // Pretrain the generator
[…]
@@
12:  // Pretrain property networks

### v2

# InstGAN: Instant Actor-Critic-Driven GAN for De Novo Molecule Generation and Property Optimization

@@ Abstract
Deep generative models, such as generative adversarial networks (GANs), have been employed for denovode~novo molecular generation in drug discovery. Most prior studies have utilized reinforcement learning (RL) algorithms, particularly Monte Carlo tree search (MCTS), to handle the discrete nature of molecular representations in GANs. However, due to the inherent instability in training GANs and RL models, along with the high computational cost associated with MCTS sampling, MCTS RL-based GANs struggle to scale to large chemical databases. To tackle these challenges, this study introduces a novel GAN based on actor-critic RL with instant and global rewards, called InstGAN, to generate molecules at the token-level with multi-property optimization. Furthermore, maximized information entropy is leveraged to alleviate the mode collapse. The experimental results demonstrate that InstGAN outperforms other baselines, achieves comparable performance to state-of-the-art models, and efficiently generates molecules with multi-property optimization. The code is available at: https://github.com/tang777777/InstGAN.
@@ 1 Introduction
Molecular graphs Shi et al. (2020) and simplified molecular input line entry systems (SMILES) strings Weininger (1988) constitute the two primary representations of molecules in deep generative models. However, generating molecules with desired chemical properties using such discrete representations is a non-trivial task. Most prior studies related to generative adversarial networks (GANs) Yu et al. (2017); Guimaraes et al. (2017); De Cao and Kipf (2018) typically update the generator by integrating the output probability of the discriminator with the chemical properties of generated molecules as a reward for reinforcement learning (RL), following the REINFORCE algorithm Williams (1992). Due to the inability of GANs to calculate rewards for partially generated molecules, Monte Carlo tree search (MCTS) is frequently utilized for sampling and completing molecules Li et al. (2022). Unfortunately, the integration of RL algorithms with GANs further exacerbates the instability of the training process. Achieving training stability with MCTS demands a substantial number of samples, rendering the process highly time-consuming Li and Yamanishi (2023).
@@
Furthermore, most aforementioned studies on denovode~novo molecular generation have focused on optimizing a single chemical property. However, in practical applications, it is often desirable to generate molecules that satisfy multiple chemical property constraints. In contrast to the former, multi-property optimization is highly complicated and challenging to achieve in nature. This is because the multi-property optimization task entails not only learning the semantic and syntactic rules of molecules to generate valid molecules from scratch but also finding pathways to optimize the distribution of chemical properties during the process Barshatski et al. (2021). For example, molecules exhibiting both drug-likeness and dopamine receptor (DRD2) activity represent only $1.6\%$ of the generated molecules Jin et al. (2019). Therefore, employing deep generative models for denovode~novo molecular generation with multi-property optimization holds significance to the drug discovery industry Barshatski and Radinsky (2021).
@@
Inspired by the previous studies in De Cao and Kipf (2018); Tang et al. (2023), this study introduces a novel GAN based on actor-critic RL with instant rewards (IR) and global rewards (GR), called InstGAN, to generate molecules at the token-level with multi-property optimization. Specifically, the generator is constructed using a long short-term memory (LSTM) that generates SMILES strings in an autoregressive manner. The discriminator quantifies each token based on SMILES substrings produced by the generator. To enhance the ability of the discriminator to quantize tokens, a bidirectional LSTM (Bi-LSTM) is chosen as the discriminator. Additionally, multi-property prediction networks with the same structure as the discriminator predict the corresponding property scores for each token as well. Subsequently, the scores of discriminator and property prediction networks, along with their global-level scores, serve as rewards for RL. To expedite the training process and facilitate its application to extensive chemical databases, an IR calculation based on actor-critic RL is proposed to update the generator. Furthermore, the maximized information entropy (MIE) is included in the generator loss function to mitigate mode collapse and enhance molecular diversity. The main contributions are
@@ 2.1 Variational Autoencoder (VAE)-based Models
Two VAE variants, Character-VAE and Grammar-VAE Kusner et al. (2017), integrate parse trees with VAEs to facilitate the generation of syntactically valid molecules. However, due to ignoring the semantic information of the molecular representations, the correlation between the training set and the generated molecules cannot be guaranteed. In contrast, Syntax-VAE Dai et al. (2018) ensures that the generated molecules are both syntactically valid and semantically meaningful. JT-VAE Jin et al. (2018) adopts a two-step VAE approach, first generating tree-structured scaffolds based on chemical substructures and then combining these outputs into complete molecules using a graph message-passing network Dai et al. (2016); Gilmer et al. (2017). Nonetheless, a major limitation of VAEs is the typically limited size of the latent space, which may restrict the capacity to produce highly realistic molecules.
@@ 2.2 Flow-based Models
GraphAF Shi et al. (2020) utilizes a flow-based autoregressive model to generate molecular graphs. GraphDF Luo et al. (2021) incorporates invertible modulo shift transforms to connect discrete latent variables with graph nodes and edges, resulting in the generation of molecular graphs. MoFlow Zang and Wang (2020) adopts a Glow-based model Kingma and Dhariwal (2018) to generate chemical bonds as graph edges and employs a graph conditional flow to subsequently generate atoms as graph nodes, followed by posthoc validity correction. GraphCNF Lippe and Gavves (2021) falls under the category of flow-based models, commonly applied in various data domains. In molecular graph generation, GraphCNF leverages flow-based techniques to address the unique challenges associated with generating molecular graphs. However, flow-based models exhibit several notable limitations: the computation of the Jacobian matrix is time-consuming, often requiring approximations, and the necessity for network invertibility imposes constraints on their representational capacity, limiting their flexibility in modeling complex, high-dimensional data distributions Zhang et al. (2021). Additionally, ensuring invertibility across all layers can lead to challenges in model training and optimization.
@@ 2.3 Diffusion-based Models
Recently, diffusion models have found application in the domain of molecular generation, where they are being utilized to tackle the intricate challenges associated with generating molecular structures that adhere to specific chemical and property constraints. DiGress Vignac et al. (2023) iteratively refines noisy graphs by adding or removing edges and adjusting categories, which results in the generation of molecular graphs with node and edge attributes suitable for classification. GDSS Jo et al. (2022) and D2L-OMP Guo et al. (2023) are the two SOTA models in the realm of molecular generation. GDSS skillfully captures the joint distribution of molecular nodes and edges, generating molecular replicas closely aligned with the training distribution. D2L-OMP generates molecules with property optimization based on a hybrid Gaussian distribution by employing diffusion on two structural levels: molecules and molecular fragments. This hybrid Gaussian distribution is then utilized in the reverse denoising process.
@@ 2.4 GAN-based Models
SeqGAN Yu et al. (2017) pioneered the incorporation of MCTS-based RL into GAN architecture, specifically designed to handle discrete text. This innovation has inspired various studies on molecular generation using GANs. MolGAN De Cao and Kipf (2018) introduces an implicit, likelihood-free discrete GAN for generating small molecular graphs. However, MolGAN faces an overfitting problem, leading to less than 5% uniqueness in the generated molecules. ORGAN Guimaraes et al. (2017) integrates domain-specific knowledge as rewards for generating SMILES strings through the MCTS-based RL algorithm. TransORGAN Li et al. (2022) leverages a transformer architecture to capture semantic information and employs variant SMILES technique for syntax rule learning. SpotGAN Li and Yamanishi (2023) adopts a first-decoder-then-encoder transformer model for generating SMILES strings from a given scaffold. However, the use of MCTS-based RL in GANs often demands a substantial number of samples for training stability, making it impractical for extensive chemical databases. EarlGAN Tang et al. (2023), while capable of generating valid molecules on large chemical databases, lacks the ability to optimize chemical properties, particularly multiple properties simultaneously.
This study introduces an actor-critic RL-driven GAN that employs IR and GR to enhance the efficiency of learning semantic and syntactic rules in SMILES strings. Furthermore, our other goal is to optimize molecule generation across diverse chemical properties.
@@ 3 Model
Figure 1 illustrates the overall architecture of InstGAN and highlights the three key substructures (i.e., the generator, discriminator, and multiple chemical property prediction networks) that are crucial for generating molecules with multi-property optimization from scratch. Formally, let $G_{\bm{\theta}}$ and $D_{\bm{\phi}}$ represent the generator and discriminator of InstGAN with parameters $\bm{\theta}$ and $\bm{\phi}$, respectively. ${\bm{S}}_{1:T}=[s_{1},\cdots,s_{T}]$ denotes a SMILES string with a sequence length of $T$. Then, the min-max optimization procedure Goodfellow et al. (2014) is implemented as follows:
$\min_{\bm{\theta}}\max_{\bm{\phi}}V(G_{\bm{\theta}},D_{\bm{\phi}})=$
$\mathbb{E}_{{\bm{S}}\sim p_{r}({\bm{S}})}[\log D_{\bm{\phi}}({\bm{S}})]+\mathbb{E}_{{\bm{S}}\sim p_{\bm{\theta}}({\bm{S}})}[\log(1-D_{\bm{\phi}}({\bm{S}}))],$ (1)
where $p_{r}(\cdot)$ and $p_{\bm{\theta}}(\cdot)$ represent the distributions of training SMILES strings and generated sequences, respectively. Following a similar approach to NLP methods de Masson d’Autume et al. (2019); Fedus et al. (2018), InstGAN utilizes an autoregressive generator and discriminator. Notably, this design allows for the allocation of dense rewards at the token-level. For detailed explanation of InstGAN’s design motivations and a full description of its generator and discriminator architectures, see Appendix A [^1] The Appendix is available at: https://github.com/tang777777/InstGAN/blob/main/Appendix.pdf.
Figure 1: The architectural overview of InstGAN comprises three main substructures. (a) The generator, featuring an LSTM, produces tokens in an autoregressive manner at each time step. MIE is employed to enhance the likelihood of sampling different tokens, smoothing the output probability distribution for the generator. (b) The discriminator, a Bi-LSTM, scores generated tokens based on both forward and backward at each time step, enabling semantic and syntactic discrimination at the token-level. Higher probabilities are assigned to likely generated tokens, while those with errors receive lower probabilities. The average of all-token probabilities assesses overall generation quality from a global perspective. (c) Multiple pretrained chemical property prediction networks, labeled Critic 11 to Critic NN, have the same structure as the discriminator. They predict various chemical properties for each token of a generated SMILES string. Similarly, the sum of stepwise scores is averaged to provide a global property score for the entire SMILES string. Finally, scores from the discriminator and critics serve as rewards in the actor-critic RL algorithm, co-updating the generator via the policy gradients.
@@ 3.1 Token-level Critics
Unlike graph-based approaches, molecular generative models relying on SMILES strings often face challenges in ensuring high validity due to the intricacies of checking valence during the autoregressive generation process. Typically, invalid SMILES strings may arise from mismatched tokens, requiring the removal or replacement of other suitable tokens to restore validity. A crucial aspect of addressing this issue lies in the meticulous assessment of SMILES strings at the token level. We therefore integrate the validity assessment directly into the generation process, by treating each token generation as an action in a token-level actor-critic RL loop: at every step, the generator (actor) proposes a candidate token, while the discriminator (critic) assess its syntactic validity and return a dense reward. This feedback system ensures the generator of high-validity token generation.
@@
In InstGAN, the generator employs an autoregressive process to produce SMILES strings, serving as inputs for the discriminator. Diverging from traditional discriminators, InstGAN performs token-level discrimination for each generated token. Specifically, with a Bi-LSTM, let $\overrightarrow{\bm{S}}_{1:t}$ and $\overleftarrow{\bm{S}}_{t:T}$ represent the forward and backward SMILES substrings, respectively. The discriminator calculates the probability $r^{D}_{t}$ that it deems the $t$-th token $\widetilde{s}_{t}$ of the SMILES string as true. The calculation is as follows:
@@
To assess the validity of the entire SMILES string, we calculate the global discriminator score as
@@
The chemical property prediction networks are replicated from the discriminator and share the same structure. In the pre-training phase, the entire SMILES string serves as input, and the networks utilize the corresponding property values of the complete SMILES string as labels for the chemical properties of its tokens at each step. In the training phase of InstGAN, these networks calculate both instant and global chemical properties of the generated SMILES string. Collaborating with the discriminator, they contribute to updating the generator using actor-critic RL-based rewards.
@@ 3.2 Instant and Global Reward Calculation
Following the actor-critic RL algorithm, the generator functions as the actor network responsible for action selection, while the discriminator, along with the property prediction networks, serves as the critic for reward calculation. However, in contrast to the traditional actor-critic RL algorithm Konda and Tsitsiklis (1999), which calculates the reward at the last time step $T$, we compute the reward for each token as
@@
where $W^{D}$ represents the weight assigned to the discriminator’s GR. Similarly, the reward of the property prediction networks can be calculated and denoted as $R^{C_{i}}_{t}$, $i\in[1,2,\cdots,N]$, and $N$ indicates the number of chemical properties to be optimized. Subsequently, the overall reward for chemical properties is represented as $R^{C}_{t}=\sum_{i=1}^{N}W_{C_{i}}R^{C_{i}}_{t}$, where $W_{C_{i}}$ denotes the weight assigned to the $i$-th chemical property prediction network. Finally, the total reward $R_{t}$ can be calculated as follows:
@@
where $\lambda$ represents a hyperparameter that balances the trade-off between the GAN and RL.
@@
2:Initialization: $G_{\bm{\theta}}$, $D_{\bm{\phi}}$, $C^{i}_{{\bm{\varphi}}_{i}},i\in[1,\cdots,N]$
@@
4: // Pre-train the discriminator
5:for $k=1\to\text{d-steps}$ do
@@
8: // Pre-train the generator
9:for $k=1\to\text{g-steps}$ do
@@
12: // Pre-train property networks